Convert the following temperatures to the Kelvin scale.
        25°C

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

25°C into K
25 + 273 = 298 K

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
To convert °C into K , we need to add 273.
25°C into K 
= 25 + 273 = 298 K
 
Hope it helps!
Thank you ✌️